OpenStudy (bossimbacon):

Complete the standard form of the equation y=10-5/2x ____x + _____y = 20 @uri @perl @sleepyhead314

OpenStudy (sleepyhead314):

is the x on the bottom or top?

OpenStudy (bossimbacon):

it's on the side

OpenStudy (sleepyhead314):

if x is on top multiply each value by 2 so you'll get 2y = 20 - 5x now get x to the other side by add 5x to both sides

OpenStudy (bossimbacon):

@sleepyhead314 i dont understand can you tell me the answer but tell me the explanation to the answer?

OpenStudy (sleepyhead314):

just do what I told you to do add 5x to both sides 2y = 20 - 5x +5x + 5x 5x + 2y = 20